# A Generalized Approach to Online Convex Optimization

### v2

@@ 4 Re-statement of Previous Result: Oblivious to fully adaptive adversary
The following theorem states a result that is well-known in the literature. For example, Hazan et al., (2016) only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary. Here we explicitly mention it as a theorem for completion.
@@
The OGD, FTRL, and OMD algorithms are deterministic and therefore have same regret bound in fully-adaptive setting as the oblivious setting.
@@ 5 Linear to convex with fully adaptive adversary
When $\mu=0$, we may also refer to ${\mathbf{F}}_{\mu}$ as the linearization of ${\mathbf{F}}$. We say ${\mathbf{F}}$ is closed under $\mu$-quadratization if ${\mathbf{F}}\supseteq{\mathbf{F}}_{\mu}$. Similarly, for $B>0$, we define ${\mathbf{F}}_{\mu}[B]$ to be the the class of functions $q$ defined above where instead we have ${\mathbf{o}}\in{\mathbb{B}}_{{\mathbf{0}}}(B)$.
@@ 6 Full information feedback to semi-bandit feedback
Any algorithm designed for semi-bandit setting may be trivially applied in the first-order full-information feedback setting. In the following theorem, we show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough. Specifically, we show that an algorithm that requires full information feedback could be converted to an algorithm that only requires semi-bandit feedback with the same regret bounds. The meta-algorithm that does this conversion is described in Algorithm 1. Next we show that $\operatorname{FTS}({\mathcal{A}})$ always performs at least as good as ${\mathcal{A}}$ when the oracle is deterministic.
@@
Algorithm 1 Full information to semi-bandit - $\operatorname{FTS}({\mathcal{A}})$
@@
where ${\mathcal{A}}^{\prime}=\operatorname{FTS}({\mathcal{A}})$ is Algorithm 1. In particular, if ${\mathbf{F}}$ is closed under $\mu$-quadratization, then
@@
See Appendix C for proof. Note that when the base algorithm ${\mathcal{A}}$ is semi-bandit, we have $\operatorname{FTS}({\mathcal{A}})={\mathcal{A}}$ and the above theorem becomes trivial.
Remark 2.
In Zhang et al., 2018a , Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback. Algorithm 3, i.e., “Improved Ader”, is the result of applying the $\operatorname{FTS}$ meta-algorithm to to Ader.
@@ 7 Lipschitz to non-Lipschitz
It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in Flaxman et al., (2005) and Lemma 1). Here we use a shrinking method described in Pedramfar et al., (2023). We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}^{\operatorname{aff}({\mathcal{K}})}({\mathbf{c}})\subseteq{\mathcal{K}}$. Then, for any shrinking parameter $0\leq\alpha<r$, we define $\hat{{\mathcal{K}}}_{\alpha}:=(1-\frac{\alpha}{r}){\mathcal{K}}+\frac{\alpha}{r}{\mathbf{c}}$. We define ${\mathbf{F}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to function class $\{f|_{\hat{{\mathcal{K}}}_{\alpha}}\mid f\in{\mathbf{F}}\}$. Given an adversary $\operatorname{Adv}$, we define $\operatorname{Adv}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the adversary constructed by restricting the output of $\operatorname{Adv}$ to the set $\hat{{\mathcal{K}}}_{\alpha}$. Recall that the the domain ${\mathcal{K}}$ of the adversary is known to the agent ${\mathcal{A}}$. For an online algorithm ${\mathcal{A}}$, we define ${\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the online algorithm resulting from restricting the domain of ${\mathcal{A}}$ to $\hat{{\mathcal{K}}}_{\alpha}$.
@@ 8 Deterministic feedback to stochastic feedback
In this section, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. In particular, there is no need for any variance reduction method, such as momentum Mokhtari et al., (2020); Chen et al., (2018); Xie et al., (2020), as long as we know that the algorithm is designed for a fully adaptive adversary.
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
See Appendix E for proof. Theorem 6.5 in Hazan et al., (2016) may be seen as a special case of the above theorem when ${\mathcal{U}}$ is a singleton and ${\mathcal{A}}$ is deterministic.
@@ 9 First order feedback to zeroth order feedback
When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient. The “smoothing trick” is a classical result by Flaxman et al., (2005) which involves averaging through spherical sampling around a given point. Here we use a variant that was introduced in Pedramfar et al., (2023) which does not require extra assumptions on the convex set ${\mathcal{K}}$.
@@
       ${\mathbf{x}}_{t}\leftarrow$ the action chosen by ${\mathcal{A}}$
       Play ${\mathbf{x}}_{t}$
@@
             Query the oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$ to get $o_{t,i}$
@@
Algorithm 2 First order to zeroth order - $\operatorname{FOTZO}({\mathcal{A}})$
@@
$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}^{\operatorname{aff}({\mathcal{K}})}({\mathbf{x}})}[f({\mathbf{z}})]={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}^{{\mathcal{L}}_{0}}({\mathbf{0}})}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
where ${\mathcal{L}}_{0}=\operatorname{aff}({\mathcal{K}})-{\mathbf{x}}$, for any ${\mathbf{x}}\in{\mathcal{K}}$ is the linear space that is a translation of the affine hull of ${\mathcal{K}}$ and ${\mathbf{v}}$ is sampled uniformly at random from the $k=\operatorname{dim}({\mathcal{L}}_{0})$-dimensional ball ${\mathbb{B}}_{1}^{{\mathcal{L}}_{0}}({\mathbf{0}})$. Thus, the function value $\hat{f}_{\delta}({\mathbf{x}})$ is obtained by “averaging” $f$ over a sliced ball of radius $\delta$ around ${\mathbf{x}}$. For a function class ${\mathbf{F}}$ over ${\mathcal{K}}$, we use $\hat{{\mathbf{F}}}_{\delta}$ to denote $\{\hat{f}_{\delta}\mid f\in{\mathbf{F}}\}$. We will drop the subscripts $\alpha$ and $\delta$ when there is no ambiguity.
@@
Let ${\mathbf{F}}$ be a function class over a convex set ${\mathcal{K}}$ and choose ${\mathbf{c}}$ and $r$ as described above and let $\delta\leq\alpha<r$. Let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and let $\hat{{\mathcal{U}}}=(1-\frac{\alpha}{r}){\mathcal{U}}+\frac{\alpha}{r}{\mathbf{c}}$. Assume ${\mathcal{A}}$ is an algorithm for online optimization with first order feedback. Then, if ${\mathcal{A}}^{\prime}=\operatorname{FOTZO}({\mathcal{A}})$ where $\operatorname{FOTZO}$ is described by Algorithm 2, we have:
@@
Under the assumptions of Theorem 6, if we further assume that ${\mathcal{A}}$ is semi-bandit, then the same regret bounds hold with ${\mathcal{A}}^{\prime}=\operatorname{STB}({\mathcal{A}})$, where $\operatorname{STB}$ is described by Algorithm 3.
@@
Pass $\hat{{\mathcal{K}}}_{\alpha}$ as the domain to ${\mathcal{A}}$
@@
       Sample ${\mathbf{v}}_{t}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
       ${\mathbf{x}}_{t}\leftarrow$ the action chosen by ${\mathcal{A}}$
       Play ${\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t}$
@@
       Let $o_{t}$ be the output of the value oracle
       Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
@@
Algorithm 3 Semi-bandit to bandit - $\operatorname{STB}({\mathcal{A}})$

### v3

@@ 4 Re-statement of Previous Result: Oblivious to fully adaptive adversary
The following theorem states a result that is well-known in the literature. For example, (?) only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary. Here we explicitly mention it as a theorem for completion.
@@
The $\mathtt{OGD}$, $\mathtt{FTRL}$, and $\mathtt{OMD}$ algorithms are deterministic and therefore have same regret bound in fully-adaptive setting as the oblivious setting.
@@ 5 Linear to convex with fully adaptive adversary
When $\mu=0$, we may also refer to ${\mathbf{F}}_{\mu}$ as the linearization of ${\mathbf{F}}$. We say ${\mathbf{F}}$ is closed under $\mu$-quadratization if ${\mathbf{F}}\supseteq{\mathbf{F}}_{\mu}$. Similarly, for $B>0$, we define ${\mathbf{Q}}_{\mu}[B]$ to be the the class of functions $q$ defined above where instead we have ${\mathbf{o}}\in{\mathbb{B}}_{B}({\mathbf{0}})$.
@@ 6 Full information feedback to semi-bandit feedback
Any algorithm designed for semi-bandit setting may be trivially applied in the first-order full-information feedback setting. In the following theorem, we show that in online convex optimization for fully adaptive adversary, semi-bandit feedback is generally enough. Specifically, we show that an algorithm that requires full information feedback could be converted to an algorithm that only requires semi-bandit feedback with the same regret bounds. The meta-algorithm that does this conversion is described in Meta-algorithm 1. Next we show that $\mathtt{FTS}({\mathcal{A}})$ always performs at least as good as ${\mathcal{A}}$ when the oracle is deterministic.
@@
Meta-algorithm 1 Full information to semi-bandit - $\mathtt{FTS}({\mathcal{A}})$
@@
where ${\mathcal{A}}^{\prime}=\mathtt{FTS}({\mathcal{A}})$ is Meta-algorithm 1. In particular, if ${\mathbf{F}}$ is closed under $\mu$-quadratization, then
@@
See Appendix C for proof. Note that when the base algorithm ${\mathcal{A}}$ is semi-bandit, we have $\mathtt{FTS}({\mathcal{A}})={\mathcal{A}}$ and the above theorem becomes trivial.
@@ 7 Lipschitz to non-Lipschitz
It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in (?) and Lemma 1). Here we use a shrinking method described in (?). We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}({\mathbf{c}})\cap\operatorname{aff}({\mathcal{K}})\subseteq{\mathcal{K}}$. Then, for any shrinking parameter $0\leq\alpha<r$, we define $\hat{{\mathcal{K}}}_{\alpha}:=(1-\frac{\alpha}{r}){\mathcal{K}}+\frac{\alpha}{r}{\mathbf{c}}$. We define ${\mathbf{F}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to function class $\{f|_{\hat{{\mathcal{K}}}_{\alpha}}\mid f\in{\mathbf{F}}\}$. Given an adversary $\operatorname{Adv}$, we define $\operatorname{Adv}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the adversary constructed by restricting the output of $\operatorname{Adv}$ to the set $\hat{{\mathcal{K}}}_{\alpha}$. Recall that the the domain ${\mathcal{K}}$ of the adversary is known to the agent ${\mathcal{A}}$. For an online algorithm ${\mathcal{A}}$, we define ${\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the online algorithm resulting from restricting the domain of ${\mathcal{A}}$ to $\hat{{\mathcal{K}}}_{\alpha}$.
@@ 8 Deterministic feedback to stochastic feedback
In this section, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. In particular, there is no need for any variance reduction method, such as momentum (?, ?, ?), as long as we know that the algorithm is designed for a fully adaptive adversary.
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
See Appendix E for proof. Theorem 6.5 in (?) may be seen as a special case of the above theorem when ${\mathcal{U}}$ is a singleton and ${\mathcal{A}}$ is deterministic.
@@ 9 First order feedback to zeroth order feedback
When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient. The “smoothing trick” is a classical idea in convex optimization (See (?)) which was first used in online convex optimization in (?). This idea involves averaging through spherical sampling around a given point. Here we use a variant that was introduced in (?) which does not require extra assumptions on the convex set ${\mathcal{K}}$.
@@
       Play ${\mathbf{x}}_{t}$, where ${\mathbf{x}}_{t}$ is the action chosen by ${\mathcal{A}}$
@@
             Let $o_{t,i}$ be the output of the query the oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$
@@
Meta-algorithm 2 First order to zeroth order - $\mathtt{FOTZO}({\mathcal{A}})$
Pass $\hat{{\mathcal{K}}}_{\alpha}$ as the domain to ${\mathcal{A}}$
$k\leftarrow\operatorname{dim}({\mathcal{L}}_{0})$
for $t=1,2,\dots,T$ do
       Sample ${\mathbf{v}}_{t}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
       Play ${\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t}$, where ${\mathbf{x}}_{t}$ is the action chosen by ${\mathcal{A}}$
       Let $f_{t}$ be the function chosen by the adversary
       Let $o_{t}$ be the output of the value oracle
       Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
end for
Meta-algorithm 3 Semi-bandit to bandit - $\mathtt{STB}({\mathcal{A}})$
@@
$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{K}})}[f({\mathbf{z}})]$
$={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
where ${\mathcal{L}}_{0}=\operatorname{aff}({\mathcal{K}})-{\mathbf{x}}$, for any ${\mathbf{x}}\in{\mathcal{K}}$ is the linear space that is a translation of the affine hull of ${\mathcal{K}}$ and ${\mathbf{v}}$ is sampled uniformly at random from the $k=\operatorname{dim}({\mathcal{L}}_{0})$-dimensional ball ${\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}$. Thus, the function value $\hat{f}_{\delta}({\mathbf{x}})$ is obtained by “averaging” $f$ over a sliced ball of radius $\delta$ around ${\mathbf{x}}$. For a function class ${\mathbf{F}}$ over ${\mathcal{K}}$, we use $\hat{{\mathbf{F}}}_{\delta}$ to denote $\{\hat{f}_{\delta}\mid f\in{\mathbf{F}}\}$. We will drop the subscripts $\alpha$ and $\delta$ when there is no ambiguity.
@@
Let ${\mathbf{F}}$ be a function class over a convex set ${\mathcal{K}}$ and choose ${\mathbf{c}}$ and $r$ as described above and let $\delta\leq\alpha<r$. Let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and let $\hat{{\mathcal{U}}}=(1-\frac{\alpha}{r}){\mathcal{U}}+\frac{\alpha}{r}{\mathbf{c}}$. Assume ${\mathcal{A}}$ is an algorithm for online optimization with first order feedback. Then, if ${\mathcal{A}}^{\prime}=\mathtt{FOTZO}({\mathcal{A}})$ where $\mathtt{FOTZO}$ is described by Meta-algorithm 2, we have:
@@
Under the assumptions of Theorem 6, if we further assume that ${\mathcal{A}}$ is semi-bandit, then the same regret bounds hold with ${\mathcal{A}}^{\prime}=\mathtt{STB}({\mathcal{A}})$, where $\mathtt{STB}$ is described by Meta-algorithm 3.
@@
Remark 2.
While $\mathtt{STB}$ turns a semi-bandit algorithm into a bandit algorithm, $\mathtt{FOTZO}$ turns a general first order algorithm into a zeroth order algorithm. However, we note that even when ${\mathcal{A}}$ is semi-bandit, $\mathtt{FOTZO}({\mathcal{A}})$ is not a bandit algorithm. Instead it is a full-information zeroth order algorithm that uses a single query per iteration, but its query is not the same as its action. This is because $\mathtt{FOTZO}({\mathcal{A}})$ plays the same point ${\mathbf{x}}_{t}$ as the base algorithm ${\mathcal{A}}$, but then adds a noise before querying according to ${\mathcal{A}}^{\text{query}}$.
Remark 3.
We have shown in the previous sections that for online convex optimization, semi-bandit feedback is enough. However, we have included the $\mathtt{FOTZO}$ meta-algorithm, which converts a full-information algorithm to another full-information algorithm, since the results in this section apply also to non-convex online optimization problems.
10 First order feedback to deterministic zeroth order feedback
When we have access to a deterministic zeroth order feedback, we may use the two point gradient estimator (?, ?) to significantly improve the results of the previous section.
Theorem 8.
Let ${\mathbf{F}}$ be an $M_{1}$-Lipschitz function class over a convex set ${\mathcal{K}}$ and choose ${\mathbf{c}}$ and $r$ as described above and let $\delta<r$. Let ${\mathcal{U}}\subseteq{\mathcal{K}}^{T}$ be a compact set and let $\hat{{\mathcal{U}}}=(1-\frac{\delta}{r}){\mathcal{U}}+\frac{\delta}{r}{\mathbf{c}}$. Assume ${\mathcal{A}}$ is an algorithm for online optimization with first order feedback. Then, if ${\mathcal{A}}^{\prime}=\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$ where $\mathtt{FOTZO\textrm{-}2P}$ is described by Meta-algorithm 2, we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D}{r}\right)\delta M_{1}T.$
Note that here we may choose $\delta=T^{-1}$ to see that the order of regret of ${\mathcal{A}}^{\prime}$ remains the same as that of ${\mathcal{A}}$. See Appendix G for proof.
Input :  Shrunk domain $\hat{{\mathcal{K}}}_{\delta}$, Linear space ${\mathcal{L}}_{0}$, smoothing parameter $\delta<r$, horizon $T$, algorithm ${\mathcal{A}}$
Pass $\hat{{\mathcal{K}}}_{\delta}$ as the domain to ${\mathcal{A}}$
@@ 10 First order feedback to deterministic zeroth order feedback
       Play ${\mathbf{x}}_{t}$, where ${\mathbf{x}}_{t}$ is the action chosen by ${\mathcal{A}}$
@@
             Sample ${\mathbf{v}}_{t,i}\in{\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}$ uniformly
             Let ${\mathbf{y}}_{t,i}$ be the query chosen by ${\mathcal{A}}^{\text{query}}$
             Query the deterministic oracle at the points ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$ and ${\mathbf{y}}_{t,i}-\delta{\mathbf{v}}_{t,i}$
             Pass $\frac{k}{2\delta}\left(f_{t}({\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i})-f_{t}({\mathbf{y}}_{t,i}-\delta{\mathbf{v}}_{t,i})\right){\mathbf{v}}_{t,i}$ as the oracle output to ${\mathcal{A}}$
@@
end for
Meta-algorithm 4 First order to zeroth order with Two Point gradient estimator - $\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$
11 Applications
In this section, we discuss applications of meta-algorithms discussed above to some specific base algorithms. Note that all the results stated here are immediate applications of our framework to existing results and no extra proof is needed.
Online Gradient Descent: Recall that $\mathtt{OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback that obtains a regret bound of $O\left(M_{1}T^{1/2}\right)$ over $M_{1}$-Lipschitz convex class ${\mathbf{F}}$ and $O\left(M_{1}^{2}\log(T)\right)$ over strongly convex $M_{1}$-Lipschitz convex class ${\mathbf{F}}^{\prime}$ (see (?)). Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$ and $\alpha=\delta=(T\log(T))^{-1/2}$, respectively, to see that
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{0}T^{3/4}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]=O\left(B_{0}^{2}T^{1/2}\log(T)^{1/2}\right).$
The first result generalizes the result of (?) to allow the feedback of the value oracle to be stochastic. The second result provides a new optimal algorithm for bandit strongly convex optimization, matching the result of (?). However, (?) obtains a regret bound of the same order for the strongly convex case while assuming access to a self-concordant barrier for the domain and also assuming that the functions are smooth. We do not require either of these assumptions here.
On the other hand, using Theorem 8 on $\mathtt{OGD}$, we see that:
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right),$
where the first result is proven in (?) and the second one is novel.
Ader and Improved Ader: In (?), Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback. Algorithm 3, i.e., “Improved Ader”, is simply the result of applying the $\mathtt{FTS}$ meta-algorithm to Ader. We may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{IA})}({\mathbf{u}})[1,T]$ $=O\left(B_{0}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{3/4}\right),$ (1)
where ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$. This matches the SOTA result of (?) for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback. Moreover, by applying Theorem 8 to Improved Ader, we recover another result of (?)
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{IA})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{1/2}\right),$
Online Gradient Descent with Separation Oracles: Similarly, the $\mathtt{SO\textrm{-}OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback obtains a weakly adaptive regret bound of $O(M_{1}T^{1/2})$ (Theorem 14 in (?)) Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(B_{0}T^{3/4}\right).$ (2)
This matches the result of Theorem 15 in (?) and generalizes it to allow for stochastic feedback. Moreover, using Theorem 8, we see that
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(M_{1}T^{1/2}\right),$ (3)
which is a novel result for adaptive regret of convex functions with deterministic zeroth order feedback.